OpenStudy (anonymous):

How do you explain this please : if 1 + 1 = 2 and why is 0 + 0 = 0 give me your definite answer please

OpenStudy (anonymous):

2 is defined as 1+1, I believe.

OpenStudy (anonymous):

but I can not find the true value of these simple math

OpenStudy (anonymous):

0 is the netreul element, so if you add it to anything it doesn't change so 5+0=5 and 0+0=0.

OpenStudy (bahrom7893):

is this a joke?

OpenStudy (anonymous):

O_O

OpenStudy (bahrom7893):

If you have two apples, you give one to Billy, how many apples will you have?

OpenStudy (anonymous):

yes Thomas you are correct if you measure it by decimal point but there is more than that

OpenStudy (anonymous):

If I have 50 cookies and I give ZERO to any one else, and they had none in the first place, then how many will they have?

OpenStudy (anonymous):

and Bahrom7893 sire this is absolutely not a joke thank you if you think so

OpenStudy (anonymous):

Emmy82650 yes that is the question how would you explain those "nothing" value ?

OpenStudy (anonymous):

It's a good question. I'm certain of about 0+0=0, not so much 1+1=2.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/0_(number)#As_a_number

OpenStudy (anonymous):

Here's how it works: You start with a set of numbers: usually the reals. Then you define an operation that takes two numbers and makes another number. So for every pair a,b you need to define a+b. for 1,1 we defined it as 2. Now that you have an operation, if there's a element b, with the property a+b=a for every a, then you also have a neutral element. We call that element 0.

OpenStudy (anonymous):

for 0 + 0 = 0 I think of some value how about 34 + 100 - 26 + 53 : 0 = 0 is this last value will always be zero "nothing" in the end whatever value you had count

OpenStudy (anonymous):

Thank you for the answers you are all Genius for some reason

OpenStudy (anonymous):

You're welcome, it can be very confusing I know.

OpenStudy (anonymous):

indeed....

OpenStudy (anonymous):

Thomas if I have 2 Apples and give someone else 1 apple that mean I have 1 apple left , and what if that one apple I have then I don't eat it right now and maybe not, so how many apple do I have ?

OpenStudy (anonymous):

If you have two apple and give one away, you have one apple. That's just the way the world works.